I am trying to solve by factoring, and I do not know if I am doing this correctly. Here is my work and answer:
If f(x) = 20x and g(x) = x - 4, what is g(g(f(3))?
f(3) = 20(3)
g(60) = 60 - 4
g(56) = 56 - 4
= 52 [this is my answer, but I need someone to check the answer and walk me through the solution.]
Thanks in advance for any assistance.

Answer by stanbon(75887)

If f(x) = 20x and g(x) = x - 4, what is g(g(f(3))?
-------
g(g(f(3))
---
= g(g(60))
---
= g(56)
-----
= 52
============
Cheers,
Stan H.